The market determined price in a perfectly competitive industry is P = Rs. 10. Suppose that the total cost equation of an individual firm in the industry is given by the expression
TC 1000+2Q+0.01Q2
a) What is the firm’s profit-maximizing output level and profit? Is this profit normal profit or supper normal profit? Justify your answer
b) At profit maximizing level what is firm total cost, total revenue and marginal cost
c) Why does a competitive firm is considered as a price taker and Monopoly firm as a price maker
1
Expert's answer
2020-10-21T10:17:42-0400

a)

"MC=TC'=2+0.02Q"

"MC=p"


"2+0.02Q=10"


"0.02Q=8"


"Q=400"


"TC(400)=1000+2\times400+0.01\times 400^2=3400"


"TR=10\times 400=4000"


"\pi=TR-TC"


"\pi=600"

This is a normal profit.


b)


"\pi=10Q-1000-2Q-0.01Q^2=8Q-1000-0.01Q^2"


"\frac {\delta \pi}{\delta Q}=8-0.02Q"


"Q=400"

So, TC=3400, TR=4000, MC=10


c) Since the marginal revenue of the monopoly is always below the demand curve, the price will always be above the marginal cost in equilibrium, providing the firm with economic profit.